# ZeroMerge: Parameter-Free KV Cache Compression for Memory-Efficient Long-Context LLMs

### v1

@@ Abstract
The linear growth of key-value (KV) cache memory and quadratic computational complexity pose significant bottlenecks for large language models (LLMs) in long-context processing. While existing KV cache optimization methods address these challenges through token pruning or feature merging, they often suffer from irreversible information loss or require costly parameter retraining. We propose ZeroMerge, a dynamic zero-shot compression framework that achieves efficient cache management through three key innovations: (1) Fine-grained memory allocation guided by multi-dimensional token importance metrics at head-level granularity, (2) A residual merging mechanism that preserves critical context through compensated attention scoring, and (3) Parameter-free adaptation compatible with diverse LLM architectures without retraining. Comprehensive evaluations across LLaMA-2 model demonstrate that ZeroMerge maintains full-cache performance at 5% compression ratios while doubling inference throughput at 40K token lengths. The method effectively balances memory efficiency, generation quality, and deployment flexibility, advancing practical long-context LLM applications. The code is available at https://github.com/SusCom-Lab/ZeroMerge.
@@ 1 Introduction
The advancement of the Transformer architecture has brought about revolutionary changes in handling sequence data, with large language models (LLMs) being a prime illustration. LLMs have achieved significant progress, surpassing humans in diverse applications. However, handling long contexts has posed challenges, particularly the linear expansion of the KV cache in proportion to the sequence length. For instance, the deployment of a Llama 2 7B model consumes 26GB of memory. When the token length reaches 32K, the 32G of memory occupied by the KV cache becomes the main memory consumer. This, along with the quadratic computational complexity, has impeded LLM integration and performance, spurring research on optimizing the KV cache for enhanced inference efficiency.
Abundant innovative studies on architecture compression alleviate the shortage of memory (Sarah et al., 2024; Ashkboos et al., 2024). Parameter quantization techniques(Dettmers et al., 2022; Xiao et al., 2023a) further reduce memory footprint by representing KV cache with low-bit precision (e.g., 4-bit integers), though at the cost of numerical accuracy preservation. More efforts focus on the management of the KV cache to directly address this challenge. Some sparse methods(Zhang et al., 2024b, a) reserve partial top-important tokens and eliminate others to mitigate the harm to the original generation quality. These selection criteria are based on the historical contribution in the decoding procedure, spatio-temporal relationship, or data distribution. One-sided and short-sighted pre-pruning strategies inevitably disrupt the distribution of KV embeddings and introduce unrecoverable information loss. Hybrid approaches combining quantization and sparsity (Dettmers et al., 2023) attempt to balance accuracy and compression ratio, yet still require delicate bit allocation strategies. Low rank compression techniques are utilized to correct the systematic drawback in some works (Dong et al., 2024). However, the newly introduced architectures or the retraining requirement of the model parameters constitutes a significant burden. One desirable KV management strategy should 1) strictly control the memory consumption to break the context length limits 2) minimize the generation quality degradation 3) be easily applied to various LLM backbones without extra refinement cost.
@@
Motivated by these goals, we proposed ZeroMerge, a more efficient and comprehensive dynamic zero-shot KV management algorithm.
  1. 1.
@@
    Fine-grained memory allocation. ZeroMerge makes full use of the historical contribution of tokens, spatio-temporal characteristics, and its own data distribution characteristics to conduct fine-grained memory management at the head level, making the method more adaptable to different types of long text tasks. It has achieved superior generation quality compared to the sparse strategy and is comparable to the retrained model on models. The performance of 5% of the KV cache bugdet is still on par with the complete model.
  2. 2.
@@
    Low-cost adaptation and deployment. ZeroMerge does not introduce any model parameters and does not require retraining or model fine-tuning to adapt to different compression ratios or various task types. It is adapted to various mainstream model architectures such as Llama, Falcon, and Mistral.
  3. 3.
    Efficient compression strategy. ZeroMerge adopts a light weight operational overhead of linear operational complexity and achieves more than twice the throughput of the original model at a token length of 50K. We have proved that the cache compression mechanism adopted by ZeroMerge does not drown out the original contribution of those intact tokens as the processing sequence grows.
@@ 2 Related work
Figure 1: Comparative analysis of KV cache optimization strategies: (a) Model architecture compression reduces both model parameters and KV cache size via multi-query attention heads; (b) Parameter/data quantization trades numerical precision for cache footprint reduction through low-bit representations; (c) Sparse approximation directly addresses memory-compute growth by evicting low-contribution tokens, albeit inducing attention distribution drift; (d) Token merge mitigates sparse caching errors by learning residuals between full and sparse attention outputs, yet introduces overhead from auxiliary networks or fine-tuning; (e) ZeroMerge (proposed) achieves zero-shot compression competitive with sparse+residual methods without added parameters, while demonstrating superior generalization across task domains and compression ratios.
The phenomenon of emergence in large-scale models has ignited new enthusiasm and provided fresh directions for exploring the nonlinear leaps in reasoning and generative capabilities of LLMs—driven by the expansion of model scale and complexity. The unprecedented parameter sizes and intensive computational demands have endowed LLMs with remarkable reasoning and generative abilities, yet they also pose significant challenges, hindering their cost-effective and large-scale deployment across various scenarios. Numerous efforts have been made to address the issue.
2.1 Model architecture compression
Works such as MiniLLM (Gu et al., 2024) and SliceGPT (Ashkboos et al., 2024) have been specifically devoted to ensure that the performance of sample tasks remains intact for more compact models. MiniLLM focuses on knowledge distillation of large language models, aiming to transfer the knowledge from large and powerful models to smaller ones, thereby reducing the number of parameters in the latter. On the other hand, SliceGPT adopts a different approach by replacing weight matrix with a smaller dense matrix. Through these innovative methods, they both strive to strike a balance between controlling the parameter count and maintaining satisfactory performance on relevant tasks. In the research of attention mechanisms, (Voita et al., 2019) pointed out that the contributions of different attention heads to the overall performance of the model in the multi-head self-attention mechanism are not equal. It also provided a novel pruning method at the attention head level. Similarly, a large number of works related to Multi-Query Attention (MQA) enable different attention “head” to share a set of “keys” and “values”, thus significantly reducing the size of these tensors and the memory bandwidth requirements for incremental decoding (Shazeer, 2019; Ainslie et al., 2023).
2.2 Parameter and data quantization
Quantization techniques have emerged as a crucial approach to address the computational and storage challenges of large language models. GPTQ(Frantar et al., 2022) is a one-shot weight quantization method based on approximate second-order information. It can quantize large GPT models with a significant number of parameters in a relatively short time and with minimal accuracy degradation. This enables models to be executed on a single GPU for generative inference and provides end-to-end inference speedups when using different GPUs. INT-FlashAttention(Chen et al., 2024) is the first INT8 quantization architecture compatible with the forward workflow of FlashAttention. It significantly improves the inference speed of FlashAttention on Ampere GPUs by implementing fully INT8 activations and general matrix-multiplication kernels. It also serves as a general token-level post-training quantization framework compatible with other data formats. They not only reduce the model size but also enhance the inference speed and manage quantization errors, making large language models more accessible and efficient for various applications.
2.3 Sparse approximation
Other studies directly focus on the KV cache, another performance bottleneck induced by long sequence tasks, striving to fundamentally address the issues of storage growth and quadratic computational complexity that accompany the increase in sequence sample length. The universality of sparse attention values (Ji et al., 2021; Deng et al., 2024) has provided a new direction for improving the computational efficiency and reducing the memory footprint of models. (Roy et al., 2021) dynamically routes the input tokens to different attention heads based on their content, enabling each attention head to focus only on the relevant parts, thus enhancing the computational efficiency while maintaining the model’s performance. (Wang et al., 2020) takes advantage of the low rank property and projects it into a low-order matrix, thereby avoiding the attention layer with quadratic computational complexity. Another token aware sparse approach, also known as eviction strategy, employ diverse token importance metrics to address memory bottlenecks. The foundational work by (Xiao et al., 2023b) identifies the attention sink phenomenon, where initial tokens disproportionately attract attention scores regardless of positional distance. Building on this insight, StreamingLLM combines a sliding temporal window with persistent retention of initial token KV states, enabling stable infinite-length generation. Similarly, LM-infinite (Han et al., 2024) introduces Lambda-shaped attention masks to differentially strengthen local (recent tokens) and global (starting tokens) semantic modeling. H2O (Zhang et al., 2024b)proposes Heavy Hitter token detection through cumulative attention score thresholds, dynamically evicting non-critical tokens during decoding. Simultaneously, PyramidKV(Zhang et al., 2024a) reveals layer-wise attention divergence - scattered patterns in lower layers gradually consolidate into instruction-critical focus in upper layers. This hierarchy informs its arithmetic-progressive cache allocation, preserving full instruction tokens while progressively reducing intermediate-layer retention.These methods demonstrate multidimensional approaches to token importance quantification. However, these eviction strategies inevitably induce attention distribution drift and error propagation, as permanently discarding low-salience tokens disrupts the intrinsic attention dynamics required for contextually coherent generation (Dao et al., 2022; Xiao et al., 2023b).
2.4 Cache merging
Token merging, initially pioneered in vision transformers (Bolya et al., 2022; Renggli et al., 2022), offers a viable pathway for KV cache optimization by fusing semantically similar tokens. The ToMe (Bolya et al., 2022) framework introduces training-free token fusion through similarity-based matching, enabling layer-wise token reduction without architectural modifications. In textual tasks, Dynamic Memory Compression (Zhang et al., 2024a) extends this concept by adaptively merging spatially proximate tokens, achieving substantial compression while retaining model efficacy through fine-tuning. Further innovations proposed in LESS (Dong et al., 2024) decouple computational bottlenecks by approximating attention mechanisms, where auxiliary networks compress low-priority tokens into compact representations with minimal impact on attention dynamics. From irreversible pruning to feature-preserving merging, from static heuristics to context-aware fusion policies, learnable mergers impose architectural complexity, demanding additional expertise in network tuning and increased computational resources, which complicates deployment.
3 Methodology
@@
Building upon the conceptual foundation of ZeroMerge outlined earlier, this section delineates the technical framework underpinning its zero-shot compression mechanism. We propose a four-component methodology consisting of adaptive budget allocation, context-sensitive contribution evaluation, residual token merging, and stabilized attention projection. The interplay of these components ensures robust generalization across compression ratios and task domains without auxiliary parameters or fine-tuning, addressing limitations of prior eviction and merging strategies.
@@ 3.1 Preliminaries
Consider a transformer-based language model with $L$ layers, where each layer contains multi-head attention mechanisms. At decoding step $T$, let $\mathbf{K}_{T}\in\mathbb{R}^{T\times d}$ and $\mathbf{V}_{T}\in\mathbb{R}^{T\times d}$ denote the Key and Value cache matrices for a target attention head, formed by concatenating token-wise vectors up to the current step:
$\mathbf{K}_{T}=[\mathbf{k}_{1},\mathbf{k}_{2},\ldots,\mathbf{k}_{T}]^{\top},\quad\mathbf{V}_{T}=[\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{T}]^{\top},$ (1)
where $\mathbf{k}_{t},\mathbf{v}_{t}\in\mathbb{R}^{d}$ represent the Key and Value vectors for the $t$-th token, and $d$ is the head dimension. For the query vector $\mathbf{q}_{T}\in\mathbb{R}^{d}$ at position $T$, the scaled dot-product attention computes output $\mathbf{o}_{i}$ via:
$a_{t}^{(T)}=\frac{\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{t}/\sqrt{d})}{\sum_{s=1}^{T}\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{s}/\sqrt{d})}.\quad\mathbf{o}^{(T)}=\sum_{t=1}^{T}a_{t}^{(T)}\mathbf{v}_{t}$ (2)
@@ 3.2 Budget Allocation
ZeroMerge strategically compresses the original $T$-length cache into a budget $B\ll T$ through tripartite allocation:
$B=B_{c}+B_{r}+B_{p},$ (3)
where $B_{c}$, $B_{r}$, and $B_{p}$ govern context preservation, residual absorption, and proximity maintenance, respectively. The compressed cache matrices aggregate three complementary components:
$\mathbf{K}_{B}=[\mathbf{K}_{c}\|\mathbf{K}_{r}\|\mathbf{K}_{p}],\quad\mathbf{V}_{B}=[\mathbf{V}_{c}\|\mathbf{V}_{r}\|\mathbf{V}_{p}],$ (4)
@@
The context component $(\mathbf{K}_{c},\mathbf{V}_{c})$ retains top-$B_{c}$ tokens ranked by contribution scores $\mathbf{s}\in\mathbb{R}^{T}$, which quantify contextual saliency. The proximity component $(\mathbf{K}_{p},\mathbf{V}_{p})$ preserves the latest $B_{p}$ tokens, capturing local attention patterns. The residual component $(\mathbf{K}_{r},\mathbf{V}_{r})$ dynamically merges $B_{r}$ historically evicted tokens through learned transformations. This configuration subsumes eviction-based methods when $B_{r}=0$, while enabling continuous interpolation between sparse and dense representations through budget reallocation
@@ 3.3 Contribution Evaluation
The context contribution metric $\mathbf{s}^{(T)}\in\mathbb{R}^{T}$ dynamically quantifies each token’s cumulative influence across decoding steps. For the $t$-th token, its score $s_{t}^{(T)}$ evolves through Geometric decay integration of attention activations:
@@
where $\lambda\in[0,1]$ is a momentum factor controlling historical preservation. This recursive formulation 1) attenuates early-stage attention signals through geometric decay, 2) maintains temporal locality by restricting updates to causally valid positions, and 3) enables online computation with $O(1)$ update cost per token.

### v2

# ZSMerge: Zero-Shot KV Cache Compression for Memory-Efficient Long-Context LLMs

@@ Abstract
The linear growth of key-value (KV) cache memory and quadratic computational complexity in attention mechanisms pose significant bottlenecks for large language models (LLMs) in long-context processing. While existing KV cache optimization methods address these challenges through token pruning or feature merging, they often incur irreversible information loss or require costly parameter retraining. To this end, we propose ZSMerge, a dynamic KV cache compression framework designed for efficient cache management, featuring three key operations: (1) fine-grained memory allocation guided by multi-dimensional token importance metrics at head-level granularity, (2) a residual merging mechanism that preserves critical context through compensated attention scoring, and (3) a zero-shot adaptation mechanism compatible with diverse LLM architectures without requiring retraining. ZSMerge significantly enhances memory efficiency and inference speed with negligible performance degradation across LLMs. When applied to LLaMA2-7B, it demonstrates a 20:1 compression ratio for key-value cache retention (reducing memory footprint to 5% of baseline) while sustaining comparable generation quality, coupled with triple throughput gains at extreme 54k-token contexts that eliminate out-of-memory failures. The code is available at https://github.com/SusCom-Lab/ZSMerge.
@@ 1 Introduction
The advancement of the Transformer architecture has revolutionized sequence data processing, with Large Language Models (LLMs) being a prime illustration. LLMs have achieved remarkable progress, surpassing human capabilities in diverse applications. While multi-turn dialogue and long-context understanding are critical scenarios, handling extended sequences remains challenging due to the linear expansion of the Key-Value (KV) cache (which stores intermediate attention keys and values during generation to avoid redundant computations) in proportion to text length. For instance, deploying a LLaMA2-7B model will consume about 26GB of GPU memory; when the token length reaches 32K, the KV cache alone occupies 32G of memory, becoming the primary memory consumption. This, along with the quadratic computational complexity, significantly restricts LLM integration and performance, thus motivating recent research efforts toward optimizing KV cache utilization for enhanced inference efficiency.
To mitigate the large memory footprint of the KV cache, various optimization strategies have emerged. Some modify the core model architecture (Sarah et al., 2024; Ashkboos et al., 2024) or employ quantization techniques for lower precision representation (Dettmers et al., 2022; Xiao et al., 2023a). Other methods directly target the KV cache bottleneck during inference using context-aware techniques. Common among these is sparse approximation or token eviction, retaining only crucial tokens while discarding others (Zhang et al., 2024b, a); however, this can disrupt KV embeddings and cause information loss. To mitigate this, alternative context-aware techniques merge or fuse information from similar tokens (Dong et al., 2024), often using low-rank compression, though these may introduce architectural overhead or require model retraining.
Ideally, an effective and practical KV cache management strategy should 1) strictly control the memory consumption to break the context length limits, 2) minimize the generation quality degradation, and 3) readily applicable across diverse LLM backbones without necessitating fine-tuning or architectural changes (zero-shot compatibility).
@@
Motivated by these goals, we propose ZSMerge, an efficient, comprehensive, and dynamic zero-shot KV management algorithm, with the following key features:
@@
  • Fine-grained memory allocation. ZSMerge makes full use of the historical contribution of tokens, spatio-temporal characteristics, and its own data distribution characteristics to conduct fine-grained, head-level memory management. This adaptability allows it to maintain superior generation quality compared to the sparse strategies and achieve performance comparable to retraining-dependent approaches, even when operating with significantly reduced memory budgets. This ensures effectiveness across diverse long-context tasks.
@@
  • Zero-shot and low-cost Integration. Designed for ease of use, ZSMerge does not introduce any model parameters and thus requires no retraining or model fine-tuning to adapt to different compression ratios or various task types. This also makes it compatible to various mainstream LLM architectures such as LLaMA, Falcon, and Mistral, thereby minimizing deployment overhead.
  • Efficient and throughput-boosting strategy. Characterized by linear operational complexity, ZSMerge’s compression mechanism is computationally lightweight. This efficiency brings significant throughput improvement, e.g., it helps achieve over triple the inference throughput compared to the original model at a token length of 54K. Furthermore, we have proved that the KV cache compressed by ZSMerge can effectively preserves the informational contribution of retained tokens, preventing signal degradation even as context length increases substantially.
@@ 2 Related Work
Figure 1: Comparative analysis of KV cache optimization strategies: (A) Sparse approximation directly addresses memory-compute growth by evicting low-contribution tokens, albeit inducing attention distribution drift. (B) Token merge mitigates sparse caching errors, often by learning residuals implicitly or explicitly, yet introduces overhead from auxiliary networks or fine-tuning. (C) ZSMerge (proposed) achieves zero-shot compression competitive with methods requiring learned components, doing so without added parameters while demonstrating superior generalization.
The escalating memory and computational demands of large language models (LLMs), particularly from KV cache growth in long-context scenarios, have spurred diverse optimization strategies. These are broadly categorized into context-agnostic and contextually adaptive approaches.
2.1 Context-Agnostic Optimization
Context-agnostic methods universally reduce resource consumption by altering the model’s structure or numerical representation, irrespective of the input sequence. Foundational techniques include Structural Compression via knowledge distillation (e.g., MiniLLM (Gu et al., 2024)) or matrix factorization (e.g., SliceGPT (Ashkboos et al., 2024)). Attention Mechanism Optimization, such as head pruning (Voita et al., 2019) and key-value sharing in Multi-Query (MQA (Shazeer, 2019)) and Grouped-Query Attention (GQA (Ainslie et al., 2023)), significantly cuts memory bandwidth. Simultaneously, Numerical Optimization tackles precision redundancy through techniques like post-training quantization (e.g., GPTQ (Frantar et al., 2022)), hardware-aligned quantization (e.g., INT-FlashAttention (Chen et al., 2024)), and exploiting low-rank properties (e.g., Linformer (Wang et al., 2020)). These strategies collectively alleviate memory bottlenecks via architectural simplification and precision reduction.
2.2 Contextually Adaptive Optimization
While context-agnostic methods offer baseline efficiency, recent research increasingly aims to directly target the KV cache—a critical bottleneck in long-sequence tasks driven by storage growth and quadratic computational complexity. The observed sparsity in attention values (Ji et al., 2021; Deng et al., 2024) motivates dynamic, context-aware strategies. Figure 1 illustrates the main categories of these adaptive approaches:
(A) Sparse Approximation (Eviction): This class directly addresses memory-compute growth by evicting low-contribution tokens, based on the premise that not all history is equally vital. Work like (Xiao et al., 2023b) identified ”attention sinks,” leading to methods like StreamingLLM (Xiao et al., 2023b) which retains initial tokens alongside a sliding window. LM-infinite (Han et al., 2024) uses specialized masks to balance local and global context. H2O (Zhang et al., 2024b) employs attention score thresholds for dynamic eviction, while PyramidKV (Zhang et al., 2024a) uses layer-wise attention patterns for progressive caching. Dynamic routing (Roy et al., 2021) also fits here. However, eviction strategies inevitably induce attention distribution drift and error propagation by permanently discarding tokens (Dao et al., 2022; Xiao et al., 2023b).
(B) Token Merging: To mitigate information loss from eviction, token merging fuses similar tokens instead of discarding them. Pioneered in vision (e.g., ToMe (Bolya et al., 2022)), it applies to text as well. Dynamic Memory Compression adaptively merges spatially proximate tokens, often requiring fine-tuning. LESS (Dong et al., 2024) uses auxiliary networks to compress tokens with minimal attention impact. While merging mitigates sparse caching errors by learning residuals (implicitly or explicitly), it often introduces overhead from auxiliary networks or fine-tuning, complicating deployment.
(C) ZSMerge (Proposed): Addressing the limitations of both prior approaches, we propose ZSMerge. It achieves zero-shot compression competitive with sparse+residual methods without added parameters, while demonstrating superior generalization across task domains and compression ratios. ZSMerge offers a parameter-free, tuning-free approach, balancing efficiency and generation performance.
3 ZSMerge Methodology
@@
Building upon the conceptual foundation of ZSMerge outlined earlier, this section delineates the technical framework underpinning its zero-shot compression mechanism. We propose a four-component methodology consisting of adaptive budget allocation, context-sensitive contribution evaluation, residual token merging, and stabilized attention projection. The interplay of these components ensures robust generalization across compression ratios and task domains without auxiliary parameters or fine-tuning, addressing limitations of prior eviction and merging strategies.
@@ 3.1 Preliminaries
Consider an $L$-layer transformer with multi-head attention mechanisms. For a target attention head at decoding step $T$, the cached Key and Value matrices are defined as:
$\mathbf{K}_{T}=[\mathbf{k}_{1},\mathbf{k}_{2},\ldots,\mathbf{k}_{T}]^{\top}\in\mathbb{R}^{T\times d},$ (1)
$\quad\mathbf{V}_{T}=[\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{T}]^{\top}\in\mathbb{R}^{T\times d},$
where $\mathbf{k}_{t},\mathbf{v}_{t}\in\mathbb{R}^{d}$ represent the Key/Value vectors for the $t$-th token. For the query vector $\mathbf{q}_{T}\in\mathbb{R}^{d}$ at position $T$, the scaled dot-product attention computes output $\mathbf{o}^{(T)}$ via:
$a_{t}^{(T)}$ $=\frac{\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{t}/\sqrt{d})}{\sum_{s=1}^{T}\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{s}/\sqrt{d})},$ (2)
$\quad\mathbf{o}^{(T)}$ $=\sum_{t=1}^{T}a_{t}^{(T)}\mathbf{v}_{t}$
@@ 3.2 Budget Allocation
ZSMerge strategically compresses the original $T$-length cache into a budget $B\ll T$ through tripartite allocation:
$B=B_{p}+B_{c}+B_{r},$ (3)
where $B_{p}$, $B_{c}$, and $B_{r}$ govern context preservation, residual absorption, and proximity maintenance, respectively. The compressed cache matrices aggregate three complementary components:
$\mathbf{K}_{B}=[\mathbf{K}_{p}\|\mathbf{K}_{c}\|\mathbf{K}_{r}],\quad\mathbf{V}_{B}=[\mathbf{V}_{p}\|\mathbf{V}_{c}\|\mathbf{V}_{r}],$ (4)
@@
The proximity component $(\mathbf{K}_{p},\mathbf{V}_{p})$ preserves the latest $B_{p}$ tokens, capturing local attention patterns. The context component $(\mathbf{K}_{c},\mathbf{V}_{c})$ retains top-$B_{c}$ tokens ranked by contribution scores $\mathbf{s}^{(T)}\in\mathbb{R}^{T}$, which quantify contextual saliency. The residual component $(\mathbf{K}_{r},\mathbf{V}_{r})$ dynamically merges $B_{r}$ historically evicted tokens through learned transformations. This configuration subsumes eviction-based methods when $B_{r}=0$, while enabling continuous interpolation between sparse and dense representations through budget reallocation.
@@ 3.3 Contribution Evaluation
The contribution scores $\mathbf{s}^{(T)}\in\mathbb{R}^{T}$ dynamically quantify each token’s cumulative influence across decoding steps. For the $t$-th token, its score $s_{t}^{(T)}$ evolves through exponential decay integration of attention activations:
@@
where the decay factor $\lambda\in[0,1]$ operates as a temporal discounting parameter analogous to reinforcement learning credit assignment, governing the exponential decay of historical attention contributions. We recommend setting it to a relatively large value to balance the contributions of long-term and short-term attention scores. Throughout this paper, $\lambda$ is configured as $0.98$ unless otherwise specified, as scores from 300 steps earlier become attenuated to approximately 0.002 through exponential decay ($0.98^{300}\approx 0.002$).